If 2x+3y=24 and 2x-3y=12 then value of xy is_____________
Found 2 solutions by Cromlix, mananth:
Answer by Cromlix(4381)

2x + 3y = 24.....1
2x - 3y = 12.....2
Add 1 to 2
4x = 36
x = 9
Substitute x = 9
into 1
2(9) + 3y = 24
18 + 3y = 24
3y = 24 - 18
3y = 6
y = 2
Therefore value of xy = 9 x 2 = 18

Answer by mananth(16949)

2.00 x + 3.00 y = 24.00 .............1
Total value
2.00 x -3.00 y = 12.00 .............2
Eliminate y
multiply (1)by 1.00
Multiply (2) by 1.00
2.00 x 3.00 y = 24.00
2.00 x -3.00 y = 12.00
Add the two equations
4.00 x = 36.00
/ 4.00
x = 9.00
plug value of x in (1)
2.00 x + 3.00 y = 24.00
18.00 + 3.00 y = 24.00
3.00 y = 24.00 -18.00
3.00 y = 6.00
y = 2.00
xy=9*2=18